# A primer on quasi-convex functions in nonlinear potential theories

### v2

@@ Acknowledgements
    We split the proof into four steps.
    Step 1. Let $\tau>1$. Then
@@
    where $B_{j}=B_{\rho_{j}}(a_{j})$, $a_{j}\in A$, $\rho_{j}=\rho(a_{j})$, satisfying for every $a_{i},a_{j}$
@@
    Consider the class $\Lambda$ of subsets $Q\subset A$ with the following properties:
    either $|a-c|>\rho(a)>\rho(C)/\tau$ or $|a-c|>\rho(C)>\rho(a)/\tau$ for all $a,c\in Q$; (i)
    whenever bA,either |ab|ρ(a) for some aQor |ab|>ρ(a)>ρ(b)/τ for all $\text{whenever $b\in A$,}\quad\begin{array}[]{l}\text{either $|a-b|\leqslant\rho(a)$ for some $a\in Q$}\\ \text{or $|a-b|>\rho(a)>\rho(b)/\tau$ for all $a\in Q$.}\end{array}$ (ii)
    By Hausdorff’s maximal principle, $\Lambda$ has a maximal member $Q$. Then the family $\{B_{\rho(a)}(a):a\in Q\}$ covers $A$. Indeed, otherwise
    $K\vcentcolon=\{a\in A:\ |a-c|>\rho(a)\ \text{for all}\ a\in Q\}\neq\emptyset$
    and thus we may choose $c\in K$ with $\tau\rho(C)>\sup_{a\in K}\rho(a)$. This implies $Q\subset Q\cup\{c\}\in\Lambda$, contrary to the maximality of $Q$. Also, since $A$ is second countable, there exists a countable subset of $Q$ which still covers $A$.
    Step 2. Let $P\subset A$. Then there exists $R\subset P$ such that
    $|a-b|>\rho(a)+\rho(b)$ for all $a,b\in R$; (iii)
    whenever $a\in P$, there exists $b\in R$ with $|a-b|\leqslant\rho(a)+\rho(b)$ and $\rho(b)>\rho(a)/\tau$. (iv)
    Consider the class $\Lambda$ of those subsets $R\subset P$ satisfying (iii) and
[…]
    Again by Hausdorff’s maximal principle, there exists a maximal member $R$ of $\Lambda$, which satisfies (iv); indeed, otherwise, arguing as in the previous step, property (v) would imply that there exists $a\in P\setminus R$ such that $R\cup\{a\}\in\Lambda$, contrary to the maximality of $R$.
    Step 3. Let $\varepsilon>0$ small and $\tau\in(1,2-\varepsilon)$ such that
@@
    Let $Q\subset A$ satisfying (i) and
    $|a-b|\leqslant\rho(a)+\rho(b)$ and $\rho(b)>\rho(a)/\tau$ for all $a,b\in Q$. (vi)
    Then there exists $N\in\mathbb{N}$ depending on $n$ and $\varepsilon$ only such that #$\#Q\leqslant N$.
@@
    $Q_{1}\vcentcolon=Q\cap\left(B_{k\rho(a)}(a)\setminus\{a\}\right),\qquad Q_{2}\vcentcolon=Q\cap B_{k\rho(a)}(a)^{\mathsf{c}}.$
@@
    and distinguish two cases. If $j=1$, since by (i) and (vi), $|b-c|>\min\{\rho(b),\rho(C)\}>\rho(a)/\tau$ and similarly $|a-c|>\rho(a)/\tau$, where $|a-b|\leqslant\kappa\rho(a)$ and $\kappa\tau-1=1-\varepsilon>0$, we obtain
    $|x-c|>|a-c|-(\rho(a)/\tau)(\kappa\tau-1)>\varepsilon|a-c|;$
    if $j=2$, since $|a-b|\leqslant\rho(a)+\rho(b)$, $|a-c|>k\rho(a)$,
    $\begin{split}|b-c|>\rho(b)\ \text{or}\ |b-c|>\rho{(C)}>\rho(b)/\tau\quad&\implies\quad|b-c|-\rho(b)>(1-\tau)\rho{(C)},\\ |a-c|>\rho{(C)}\ \text{or}\ |a-c|>\rho(a)>\rho{(C)}/\tau\quad&\implies\quad|a-c|>\rho{(C)}/\tau,\end{split}$
    we obtain
    $\begin{split}|x-c|&\geqslant|a-c|+|b-c|-\rho(a)-\rho(b)>|a-c|-\rho(a)-\rho{(C)}(\tau-1)\\ &=|a-c|\left(1-\frac{\rho(a)}{|a-c|}-\frac{\rho{(C)}}{|a-c|}(\tau-1)\right)>|a-c|\left(1-\frac{1}{k}-\tau(\tau-1)\right)\\ &>\varepsilon|a-c|,\end{split}$
@@
    Therefore we have
    $\frac{|x-c|}{|a-c|}\geqslant\varepsilon$
    for all $a,x,c$ as above. At this point by (B.27) one sees that
    $\frac{|x-c|}{|a-c|}=\left|\frac{b-a}{|b-a|}-\frac{c-a}{|c-a|}\right|=|\pi(b)-\pi{(C)}|,$
    where $\pi(b),\pi{(C)}$ denote the projections of $b,c$, respectively, onto the sphere $\partial B_{1}(a)$. This implies that there is a one-to-one correspondence between points in $Q_{j}$ and their projections and #$\#Q_{j}\leqslant\kappa$ for some $\kappa\in\mathbb{N}$. Indeed, we know that $\pi(Q_{j})$ is finite, because otherwise by the compactness of $\partial B_{1}(a)$ it would have a limit point and thus there would exist points $\pi(b),\pi{(C)}$, with $|\pi(b)-\pi{(C)}|<\varepsilon$. It is clear that this argument is independent of $a$ and holds for every $\varepsilon$-distanced subset of the unit sphere $\mathbb{S}^{n-1}$, hence $\kappa$ depends only on $n$ and $\varepsilon$.
    Putting all of this together, we have #$\#Q=\#(Q_{1}\cup Q_{2}\cup\{a\})=\#Q_{1}+\#Q_{2}+1\leqslant 2\kappa+1=:N$.
    Step 4. Let $\tau$ be as in step 3 with $\varepsilon=\frac{1}{3}$ and $P\vcentcolon=\{B_{j}\}_{j\in\mathbb{N}}$ be as in step 1. Then $P$ is the union of $N$ families of pairwise disjoint balls.
    Applying step 2, we define by induction subsets $P_{k}$, $R_{k}$ of $P$, starting with $P_{0}\vcentcolon=P$ and $R_{0}\vcentcolon=\emptyset$ and letting, for $k\geqslant 1$, $P_{k}\vcentcolon=P_{k-1}\setminus R_{k-1}$ and $R_{k}\subset P_{k}$ satisfying (iii) and (iv). Note that condition (iii) implies that the family $\mathcal{B}_{k}\vcentcolon=\{B_{j}^{k}\vcentcolon=B_{\rho_{j}^{k}}(a_{j}^{k}):\ a_{j}^{k}\in R_{k}\}$ is disjointed for each $k$ and by definition
@@
    Hence we conclude the proof if we show that $P_{N+1}=\emptyset$. In fact, if $a\in P_{N+1}$, one uses property (iv) to select, for all $k=1,\dots,N$, some $b_{j}^{k}\in R_{k}$ with $|a-b_{j}^{k}|\leqslant\rho(a)+\rho_{j}^{k}$ and $\rho_{j}^{k}>\rho(a)/\tau$. Therefore $Q\vcentcolon=\{a,b_{1},\dots,b_{N}\}$ satisfies (i) and (iv), and also #$\#Q=N+1$ since the $R_{j}$’s are disjointed by definition; but this contradicts what we proved in step 3. ∎
    Next, one can globalize the local measure estimate of Lemma LABEL:app:ball by way of the Besicovitch covering Lemma LABEL:app:bes.
@@
    Let $\Omega\subset{\mathbb{R}}^{n}$ open, $G\colon\Omega\to{{\mathbb{R}}^{n}}$ and $A\subset\Omega$ with $m(A)$ finite. Suppose $G$ is differentiable at every $a\in A$ and there exists a constant $M$ such that $|\det DG|\leqslant M$ in $A$. Then
@@
    Fix $U\supset A$ open such that $m(U)\leqslant 2m(A)$; given $\varepsilon>0$, by Lemma LABEL:app:ball, for each $a\in A$ there exists some ball $B_{\rho(a)}(a)$ such that $m(G(B_{\rho(a)}(a)))\leqslant\left(M+\varepsilon\right)\cdot m(B_{\rho(a)}(a))$. Also, we may choose $\rho(a)$ in such a way that $B_{\rho(a)}(a)\subset U\cap\Omega$ for all $a\in A$, hence the map $\rho\,\colon A\to{\mathbb{R}}_{+}$ is bounded. Besicovitch’s covering theorem (Lemma LABEL:app:bes) tells us that there exists a countable subset $\{a_{i}\}_{i\in\mathbb{N}}\subset A$ such that, letting $\rho_{i}\vcentcolon=\rho(a_{i})$, $\{B_{i}\vcentcolon=B_{\rho_{i}}(a_{i})\}_{i\in\mathbb{N}}$ covers $A$ and every point $x\in{\mathbb{R}}^{n}$ is in at most $N$ of the balls $B_{\rho_{i}}(a_{i})$. Therefore $\sum_{i\in\mathbb{N}}\chi_{B_{i}}\leqslant N\chi_{U}$ and we have
@@
    $\begin{split}m(G(A))\leqslant m\!\left(G\Big{(}\bigcup_{i\in\mathbb{N}}B_{i}\Big{)}\right)\leqslant\sum_{i\in N}m(G(B_{i}))\leqslant(M+\varepsilon)\!\sum_{i\in\mathbb{N}}m(B_{i})\leqslant 2N(M+\varepsilon)m(A).\end{split}$
@@
    Let $\Omega\subset{\mathbb{R}}^{n}$ open and $G\colon\Omega\to{\mathbb{R}}^{n}$; set
    $\mathcal{S}_{G}\vcentcolon=\{x\in\Omega:\ G\ \text{is differentiable in $x$ and}\ |\det DG(x)|=0\}.$
@@
    If $m(\mathcal{S}_{G})<\infty$, then the conclusion follows immediately from Proposition B.13 with $A=\mathcal{S_{G}}$ and $M=0$. If $m(\mathcal{S}_{G})=\infty$, then consider the decomposition $\mathcal{S}_{G}=\bigcup_{k\in N}A_{k}$ with $A_{k}:=\mathcal{S}_{G}\cap B_{k}(0)$. One has $m(A_{k})=0$ for each $k$ by the previous step, and the conclusion follows from the subadditivity of $m$. ∎
@@
    For the Lipschitz function $G$, one needs to show that $m(G(k_{G}))=0$. The set of critical points can be decomposed into $k_{G}=\mathcal{N}_{G}\cup\mathcal{S}_{G}$, where $\mathcal{N}_{G}$ denotes the set of points at which $G$ is not differentiable and $\mathcal{S}_{G}$ is the set of points of differentiability where the Jacobian $DG$ is singular. Since $G$ is Lipschitz, $\mathcal{N}_{G}$ has (Lebesgue) measure zero by Rademacher’s Theorem 1.17 and hence $G(\mathcal{N}_{G})$ has measure zero as well since $G$ is Lipschitz. By the above corollary, $G(\mathcal{S}_{G})$ has measure zero and we are done. ∎

### v3

@@ Appendix B Proof of Sard’s theorem for Lipschitz functions
We split the proof into four claims.
Claim 1. Let $\tau>1$. Then
@@
where $B_{j}=B_{\rho_{j}}(a_{j})$ with $a_{j}\in A$ and $\rho_{j}=\rho(a_{j})$ satisfying
@@
Consider the class $\Lambda$ of subsets $Q\subset A$ with the following two properties:
  1. (i)
    either $|a-c|>\rho(a)>\rho(c)/\tau$ or $|a-c|>\rho(c)>\rho(a)/\tau$ for all $a,c\in Q$;
  2. (ii)
    whenever $b\in A$, either $|a-b|\leqslant\rho(a)$ for some $a\in Q$ or $|a-b|>\rho(a)>\rho(b)/\tau$ for all $a\in Q$.
By Hausdorff’s maximal principle, $\Lambda$ has a maximal member $\mathfrak{Q}$. Then the family $\{B_{\rho(a)}(a):a\in\mathfrak{Q}\}$ covers $A$. Indeed, otherwise
$K\vcentcolon=\{a\in A:\ |a-c|>\rho(a)\ \text{for all}\ a\in\mathfrak{Q}\}\neq\emptyset$
and thus we may choose $c\in K$ with $\tau\rho(c)>\sup_{a\in K}\rho(a)$. This implies $\mathfrak{Q}\subset\mathfrak{Q}\cup\{c\}\in\Lambda$, contrary to the maximality of $\mathfrak{Q}$. Also, since $A$ is second countable, there exists a countable subset of $\mathfrak{Q}$ which still covers $A$.
Claim 2. Let $P\subset A$. Then there exists $R\subset P$ such that the following two properties hold:
  1. (1)
    $|a-b|>\rho(a)+\rho(b)$ for all $a,b\in R$;
  2. (2)
    whenever $a\in P$, there exists $b\in R$ with $|a-b|\leqslant\rho(a)+\rho(b)$ and $\rho(b)>\rho(a)/\tau$.
Consider the class $\Lambda$ of those subsets $R\subset P$ satisfying (iii) and the following:
  1. (1)
    whenever $a\in P$, either $|a-b|>\rho(a)+\rho(b)$ for all $b\in R$ or (iv) holds.
Again by Hausdorff’s maximal principle, there exists a maximal member $\mathfrak{R}$ of $\Lambda$, which satisfies (iv); indeed, otherwise, arguing as in the previous step, property (v) would imply that there exists $a\in P\setminus\mathfrak{R}$ such that $\mathfrak{R}\cup\{a\}\in\Lambda$, contrary to the maximality of $\mathfrak{R}$.
Claim 3. Let $\varepsilon\in(0,1)$ and $\tau\in(1,2-\varepsilon)$ such that
@@
Let $Q\subset A$ satisfy (i) as well as the following property:
  1. (1)
    $|a-b|\leqslant\rho(a)+\rho(b)$ and $\rho(b)>\rho(a)/\tau$ for all $a,b\in Q$.
Then there exists $N\in\mathbb{N}$, depending only on $n$ and $\varepsilon$, such that #$\#Q\leqslant N$.
@@
$Q_{1}\vcentcolon=Q\cap\left(B_{\kappa\rho(a)}(a)\setminus\{a\}\right),\qquad Q_{2}\vcentcolon=Q\cap B_{\kappa\rho(a)}(a)^{\mathsf{c}}.$
@@
and distinguish two cases. If $j=1$, since by properties (i) and (vi),
$|b-c|>\min\{\rho(b),\rho(c)\}>\rho(a)/\tau$
and similarly $|a-c|>\rho(a)/\tau$, while $|a-b|\leqslant\kappa\rho(a)$, with $\kappa\tau-1=1-\varepsilon>0$, we obtain
$|x-c|>|a-c|-(\rho(a)/\tau)(\kappa\tau-1)>\varepsilon|a-c|.$
If $j=2$, note that we have the following inequalities:
$|a-b|\leqslant\rho(a)+\rho(b),\quad|a-c|>\max\{\kappa\rho(a),\rho(c)/\tau\},\quad|b-c|-\rho(b)>(1-\tau)\rho{(c)};$
then we obtain
$\begin{split}|x-c|&\geqslant|a-c|+|b-c|-\rho(a)-\rho(b)>|a-c|-\rho(a)-(\tau-1)\rho{(c)}\\ &=|a-c|\biggl{(}1-\frac{\rho(a)}{|a-c|}-(\tau-1)\frac{\rho{(c)}}{|a-c|}\biggr{)}>|a-c|\biggl{(}1-\frac{1}{\kappa}-\tau(\tau-1)\biggr{)}\\ &>\varepsilon|a-c|,\end{split}$
@@
At this point by (B.2) one sees that
$\varepsilon<\frac{|x-c|}{|a-c|}=\left|\frac{b-a}{|b-a|}-\frac{c-a}{|c-a|}\right|=|\pi(b)-\pi{(c)}|,$
where $\pi(b)$ and $\pi{(c)}$ denote the projections of $b$ and $c$, respectively, onto the sphere $\partial B_{1}(a)$. This implies that there is a one-to-one correspondence between points in $Q_{j}$ and their projections and #$\#Q_{j}\leqslant K$ for some $K\in\mathbb{N}$. Indeed, we know that $\pi(Q_{j})$ is finite, because otherwise by the compactness of $\partial B_{1}(a)$ it would have a limit point and thus there would exist points $\pi(b)$ and $\pi{(c)}$, with $|\pi(b)-\pi{(c)}|<\varepsilon$. Clearly this argument is independent of $a$ and holds for every $\varepsilon$-distanced subset of the unit sphere $\mathbb{S}^{n-1}$, hence $K$ depends only on $n$ and $\varepsilon$. Putting all of this together, we have #$\#Q=\#(Q_{1}\cup Q_{2}\cup\{a\})=\#Q_{1}+\#Q_{2}+1\leqslant 2K+1=\vcentcolon N$.
Claim 4. Let $\tau$ be as in Claim 3 with $\varepsilon=\frac{1}{3}$ and $P\vcentcolon=\{B_{j}\}_{j\in\mathbb{N}}$ be as in Claim 1. Then $P$ is the union of $N$ families of pairwise disjoint balls.
Applying Claim 2, we define by induction subsets $P_{k},R_{k}\subset P$, starting with $P_{0}\vcentcolon=P$ and $R_{0}\vcentcolon=\emptyset$ and letting, for $k\geqslant 1$, $P_{k}\vcentcolon=P_{k-1}\setminus R_{k-1}$, and $R_{k}\subset P_{k}$ satisfying (iii) and (iv). Note that condition (iii) implies that the family $\mathcal{B}_{k}\vcentcolon=\{B_{j}^{k}\vcentcolon=B_{\rho_{j}^{k}}(a_{j}^{k}):\ a_{j}^{k}\in R_{k}\}$ is disjointed for each $k$ and by definition
@@
Hence we conclude the proof if we show that $P_{N+1}=\emptyset$. In fact, if $a\in P_{N+1}$, one uses property (iv) to select, for all $k=1,\dots,N$, some $b_{j}^{k}\in R_{k}$ with $|a-b_{j}^{k}|\leqslant\rho(a)+\rho_{j}^{k}$ and $\rho_{j}^{k}>\rho(a)/\tau$. Therefore $Q\vcentcolon=\{a,b_{1},\dots,b_{N}\}$ satisfies (i) and (iv), and also #$\#Q=N+1$ since the $R_{j}$’s are disjointed by definition; but this contradicts Claim 3, so it must be $P_{N+1}=\emptyset$ and the proof is concluded. ∎
Next, one can globalize the local measure estimate of Lemma B.1 by way of the Besicovitch covering Lemma B.2.
@@
Let $\Omega\subset{\mathbb{R}}^{n}$ be open, $G\colon\Omega\to{{\mathbb{R}}^{n}}$ and let $A\subset\Omega$ have $m(A)$ finite. Suppose $G$ is differentiable at every $a\in A$ and there exists a constant $M$ such that $\lvert\det DG\rvert\leqslant M$ in $A$. Then
@@
Fix $U\supset A$ open such that $m(U)\leqslant 2m(A)$; given $\varepsilon>0$, by Lemma B.1, for each $a\in A$ there exists some ball $B_{\rho(a)}(a)$ such that
$m(G(B_{\rho(a)}(a)))\leqslant(M+\varepsilon)m(B_{\rho(a)}(a)).$
Also, we may choose $\rho(a)$ in such a way that $B_{\rho(a)}(a)\subset U\cap\Omega$ for all $a\in A$; hence the map $\rho\,\colon A\to{\mathbb{R}}_{+}$ is bounded. Besicovitch’s covering theorem (Lemma B.2) tells us that there exists a countable subset $\{a_{i}\}_{i\in\mathbb{N}}\subset A$ such that, letting $\rho_{i}\vcentcolon=\rho(a_{i})$ and $B_{i}\vcentcolon=B_{\rho_{i}}(a_{i})$, the family $\{B_{i}\}_{i\in\mathbb{N}}$ covers $A$ and every point $x\in{\mathbb{R}}^{n}$ is in at most $N$ of the balls $B_{i}$. Therefore $\sum_{i\in\mathbb{N}}\chi_{B_{i}}\leqslant N\chi_{U}$ and we have
@@
$\begin{split}m(G(A))\leqslant m\biggl{(}G\biggl{(}\,\bigcup_{i\in\mathbb{N}}B_{i}\biggr{)}\biggr{)}\leqslant\sum_{i\in N}m(G(B_{i}))\leqslant(M+\varepsilon)\!\sum_{i\in\mathbb{N}}m(B_{i})\leqslant 2N(M+\varepsilon)m(A).\end{split}$
@@
Let $\Omega\subset{\mathbb{R}}^{n}$ be open and $G\colon\Omega\to{\mathbb{R}}^{n}$; set
$\mathcal{S}_{G}\vcentcolon=\{x\in\Omega:\ G\ \text{is differentiable at $x$ and}\ \lvert\det DG(x)\rvert=0\}.$
@@
If $m(\mathcal{S}_{G})<\infty$, then the conclusion follows immediately from Proposition B.3 with $A=\mathcal{S}_{G}$ and $M=0$. If $m(\mathcal{S}_{G})=\infty$, then consider the decomposition $\mathcal{S}_{G}=\bigcup_{k\in N}A_{k}$ with $A_{k}\vcentcolon=\mathcal{S}_{G}\cap B_{k}(0)$. One has $m(A_{k})=0$ for each $k$ by the previous step, and the conclusion follows from the subadditivity of $m$. ∎
@@
One needs to show that $m(G(k_{G}))=0$. The set of critical points can be decomposed into $k_{G}=\mathcal{N}_{G}\cup\mathcal{S}_{G}$, where $\mathcal{N}_{G}$ denotes the set of points at which $G$ is not differentiable and $\mathcal{S}_{G}$ is the set of points of differentiability where the Jacobian $DG$ is singular. Since $G$ is Lipschitz, $\mathcal{N}_{G}$ has measure zero by Rademacher’s Theorem 1.19 and hence $G(\mathcal{N}_{G})$ has measure zero as well since $G$ is Lipschitz. By the above corollary, $G(\mathcal{S}_{G})$ has measure zero and we are done. ∎